8. Suppose you tossed a coin 100 times and got 77 heads and 23 tails. Does this seem like a rea- sonable result? What inference might you draw from the result?
Answer #1

here mean number of heads =np=100*0.5=50

and std deviation =(np(1-p))1/2 =(100*0.5*(1-0.5))1/2 =5

as 77 falls more than 3 standrd deviaiton from mean ; it is a highly unusual number of heads considering coin is fair.

therefore we can infer that coin is biased towards head,